40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Thu Mar 28 03:37:07 2013, Zach, Optics, Configuration, Test setup input optics progress OMC_test_setup_IO.jpg2013-03-27_05.48.20.jpg2013-03-28_02.29.36.jpg
    Reply  Fri Mar 29 08:55:00 2013, Zach, Optics, Configuration, Beam launched into fiber 
       Reply  Thu Apr 4 23:44:52 2013, Koji, Optics, Configuration, Beam launched into fiber 
          Reply  Fri Apr 5 18:18:36 2013, Zach, Optics, Configuration, AOM probably broken Isomet_1250C_zin_4_5_13.png
Message ID: 87     Entry time: Fri Mar 29 08:55:00 2013     In reply to: 86     Reply to this: 97
Author: Zach 
Type: Optics 
Category: Configuration 
Subject: Beam launched into fiber 

Quote:

Lisa and I concurred that it felt like we had converged to the optimum alignment and polarization, which would mean that the lack of coupling is all from mode mismatch. Since the input mode is well collimated, it seems unlikely that we could be off enough to only get ~1% coupling. One possibility is that the collimator is not well attached to the fiber itself. Since the Rayleigh range within it is very small, any looseness here can be critical.

My hypothesis about the input-side collimator turned out to be correct.

I removed the fiber from the collimator and mount at the input side, and then injected the illuminator beam from this side. Since we already saw a nice (but dim) IR beam emerging from the output side the other night, it followed that that collimator was correctly attached. With the illuminator injected from the input side, I also saw a nice, collimated red beam emerging from the output. So, the input collimator was not properly attached during our previous attempts, leading to the abysmal coupling.

The problem is that the mount does not allow you to remove and reattach the fiber while the collimator is already attached, and the dimensions make it hard to fit your fingers in to tighten the fiber to the collimator once the collimator is in the mount. I disassembled the mount and found a way to attach/reattach the fiber that preserves the tight collimator contact. I will upload a how-to shortly.

With this fix, I was able to align the input beam and get decent coupling:

EOM path: ~70%

AOM path: ~50%

ELOG V3.1.3-